# Fully Spiking Neural Networks for Unified Frame-Event Object Tracking

### v1

@@ Abstract
The integration of image and event streams offers a promising approach for achieving robust visual object tracking in complex environments. However, current fusion methods achieve high performance at the cost of significant computational overhead and struggle to efficiently extract the sparse, asynchronous information from event streams, failing to leverage the energy-efficient advantages of event-driven spiking paradigms. To address this challenge, we propose the first fully Spiking Frame-Event Tracking framework called SpikeFET. This network achieves synergistic integration of convolutional local feature extraction and Transformer-based global modeling within the spiking paradigm, effectively fusing frame and event data. To overcome the degradation of translation invariance caused by convolutional padding, we introduce a Random Patchwork Module (RPM) that eliminates positional bias through randomized spatial reorganization and learnable type encoding while preserving residual structures. Furthermore, we propose a Spatial-Temporal Regularization (STR) strategy that overcomes similarity metric degradation from asymmetric features by enforcing spatio-temporal consistency among temporal template features in latent space. Extensive experiments across multiple benchmarks demonstrate that the proposed framework achieves superior tracking accuracy over existing methods while significantly reducing power consumption, attaining an optimal balance between performance and efficiency. The code will be released.
@@ 3.1 Spiking Frame-Event Tracking Network
U=U+SepSpikeConv($\textbf{U})$ = U + SepSpikeConv ( U ) (1)
@@
U=U+SepSpikeConv($\textbf{U})$ = U + SepSpikeConv ( U ) (6)
@@
where $\text{Linear}(\cdot)$ denotes a linear transformation, $\gamma$ is the channel expansion factor (zhang2023rethinking, ), $\text{CSWinSSA}(\cdot)$ represents the CSWinSSA operator, and $\text{SSA}(\cdot)$ is the spiking self-attention operator within the CSWinSSA operator. To address the challenge of large values generated by matrix multiplication, a scaling factor (scale) is introduced to regulate the magnitude of the results. Further details of the CSWinSSA operator are provided in the Appendix.
@@
In object tracking tasks, the response maps output by the network play a critical role in target localization. To achieve sufficient fusion of image and event modalities, we propose enhancing feature alignment by constraining the similarity between the response maps of the two modalities, implemented via a weighted focal loss function $\mathcal{L}_{\text{GWF}}$ (law2018cornernet, ) (see the Appendix for more details). The final loss function is expressed as: $\mathcal{L}_{\text{Res}}=\mathcal{L}_{\text{GWF}}\left(\mathbf{R}_{F}/\tau,\mathbf{R}_{E}/\tau\right)$, where $\mathbf{R}_{F}$ and $\mathbf{R}_{E}$ are the response maps of two modes, $\tau$ is the temperature coefficient, and the empirical setting is 2.
@@ 3.4 Training and Inference
where $\lambda_{\text{iou}}=2$ and $\lambda_{\text{L1}}=5$ are regularization parameters as in (yan2021learning, ). $\alpha=1$ and $\beta=0.5$ are hyperparameters used to balance the contributions of different loss functions.
@@
where $\lambda=0.5$ is a hyperparameter to balance the responses of different modalities. Notably, we do not employ a dynamic template update strategy and follow common practice by applying a Hanning window penalty to leverage positional priors in tracking (chen2021transformer, ; ye2022joint, ).
@@ 4.1 Implementation details
Our proposed SpikeFET was implemented with PyTorch 1.12 in Python 3.8 and trained on two NVIDIA RTX 4090 GPUs. The network input consists of a triplet image group comprising one search image and two distinct template images. When constructing training samples from randomly sampled video sequences, we expanded the search regions and template regions to 4× and 2× their original bounding box sizes, respectively, followed by resizing them to 256×256 and 128×128. We follow the standard configuration pipeline of OSTrack (ye2022joint, ). Specifically, we adapt our SpikeFET into a single-modal SpikeFT model, pre-trained on benchmark datasets such as COCO (lin2014microsoft, ), LaSOT (fan2019lasot, ), TrackingNet (muller2018trackingnet, ), and GOT-10K (huang2019got, ) to fine-tune our SpikeFET network. For the single-modal SpikeFT model, we remove one feature extraction branch and the spiking tracking head branch from SpikeFET, along with eliminating modality encoding. To balance speed and accuracy, we propose three variants with distinct architectures and parameters: SpikeFET-Base (Fig. 3) with larger parameter capacity, and SpikeFET-Tiny and SpikeFT optimized for computational efficiency. Please refer to the Appendix for more details. To quantitatively evaluate the performance of tracking systems, we employ three core metrics: Area Under the Curve (AUC), Precision Rate (PR), and theoretical power consumption. For a detailed description of the theoretical power consumption calculation method, please refer to Appendix.
@@ 4.2 Main Results
We evaluated the performance of our proposed SpikeFET-Tiny and SpikeFET-Base on several benchmarks, including FE108 (zhang2021object, ), VisEvent (wang2023visevent, ), and COESOT (tang2022revisiting, ). Meanwhile, for comparison with event-based trackers, we constructed a baseline model named SpikeET that using only event streams as input. Detailed dataset specifications are detailed in the Appendix.
@@
Table 1: Comparison with state-of-the-art trackers on FE108 (zhang2021object, ), VisEvent (wang2023visevent, ), and COESOT (tang2022revisiting, ). We train and measure energy consumption on the VisEvent dataset. The best three results are shown in red, blue and green fonts.
@@
ANN DiMP50 (bhat2019learning, ) - - 57.1 85.1 47.8 67.0 58.9 67.1
PrDiMP50 (danelljan2020probabilistic, ) - - 59.0 87.7 47.6 65.3 57.9 69.6
SiamRCNN (voigtlaender2020siam, ) - - - - 49.9 65.9 60.9 71.0
TrDiMP50 (wang2021transformer, ) - - 60.3 91.2 - - 60.1 72.2
TransT50 (chen2021transformer, ) - - 63.9 93.0 47.4 65.0 60.5 72.4
ToMP101 (mayer2022transforming, ) - - 61.8 91.1 - - 59.9 67.2
FENet (zhang2021object, ) - - 63.1 91.8 - - - -
@@
SpikeFET-Base 105.48 102.61 68.7 97.0 59.0 75.3 68.5 81.7
As shown in Tab. 1, compared to several state-of-the-art trackers, our SpikeFET demonstrates significant advantages across three datasets. On the FE108 (zhang2021object, ) dataset, SpikeFET-Tiny achieves an AUC score of 68.5%, matching the previous state-of-the-art HRMonTrack (zhu2023cross, ) while surpassing it by 0.3% in PR score and outperforming all other trackers by a substantial margin. Compared to the Tiny variant, the Base model further improves performance metrics, but there is a tendency towards overfitting. Notably, although the model size of the base model is comparable to the previous Base models, its theoretical energy consumption is significantly lower than equivalently scaled models. Additionally Compared with the latest SDSTrack (hou2024sdstrack, ) on ANN, our Tiny has improved its AUC score by 2.7% on the FE108 (zhang2021object, ) dataset, with parameters only one-third of SDSTrack (hou2024sdstrack, ), while reducing power consumption by 39 times.
On the VisEvent (wang2023visevent, ) dataset, our Tiny model outperforms the majority of trackers. However, for the Base model, compared with the frame-event tracker ViPT (zhu2023visual, ) based on prompt learning, our AUC shows a 0.2% decrease. We attribute this primarily to suboptimal performance in SNNs pre-training and the relatively fewer extreme scenarios in the VisEvent (wang2023visevent, ) dataset compared to FE108 (zhang2021object, ). These factors limit SNNs from fully leveraging the advantages of event data, thereby reducing their effectiveness. This observation highlights that our SNN-based network still needs to bridge the performance gap with state-of-the-art ANNs architectures in tracking applications, though the energy efficiency advantages of SNNs remain highly competitive.
@@
ANN DiMP50 (bhat2019learning, ) - - 57.1 85.1 31.5 44.2
PrDiMP50 (danelljan2020probabilistic, ) - - - - 32.2 46.9
ATOM (danelljan2019atom, ) - - - - 28.6 47.4
SiamRPN (li2018high, ) - - - - 24.7 38.4
@@ 4.3 Ablation Study
We analyze the effects of the proposed RPM, $\text{E}_{\text{t}}$, and STR. Our baseline model, similar to CEUTrack (tang2022revisiting, ), processes six input images from both frame and event modalities with separate feature extraction. These features are concatenated before entering the fusion network, then separated prior to the prediction head and processed through decoupled tracking heads. As shown in Tab. 4, the baseline equipped with RPM (#1) achieves significant improvements. On FE108 (zhang2021object, ) with numerous extreme scenarios, RPM substantially increases AUC by 34.7% and PR by 49.9% over the baseline, while on VisEvent (wang2023visevent, ), it improves AUC by 10.4% and PR by 11.9%. More specifically, illustrated in Fig. 6, without RPM method, the peak response in the output image fails to precisely indicate the target position, exhibiting a significant deviation. In contrast, when RPM is employed, the target area is accurately highlighted with high response activation. Quantitative analysis confirms that our model achieves enhanced tracking accuracy and robustness through RPM. Further additions of $\text{E}_{\text{t}}$ (#2) and STR (#3) to #1 demonstrate effective enhancements, particularly with STR improving FE108 (zhang2021object, ) AUC by 0.6%. This proves that while retaining the advantage of RPM in dynamically reorganizing spatial distributions, the STR strategy effectively suppresses matching deviations caused by feature spatial misalignment, significantly improving model performance. Finally, integrating both $\text{E}_{\text{t}}$ and STR with RPM (#4) yields greater gains: 1.7% AUC and 1.5% PR improvements on FE108 (zhang2021object, ), along with 1.3% AUC and 1.1% PR improvements on VisEvent (wang2023visevent, ).
Table 3: Ablation studies on different components on the FE108 (zhang2021object, ) and VisEvent (wang2023visevent, ) testing sets. RPM, $\text{E}_{\text{t}}$ and STR denote Random Patchwork Module, Type Encodings and Spatial-Temporal Regularization.
# RPM $\text{E}_{\text{t}}$ STR FE108 VisEvent
@@
Baseline $\times$ $\times$ $\times$ 32.1 45.0 45.1 60.5
1 $\checkmark$ $\times$ $\times$ 66.8 94.9 55.5 72.4
2 $\checkmark$ $\checkmark$ $\times$ 66.8 95.0 55.7 72.3
3 $\checkmark$ $\times$ $\checkmark$ 67.4 95.8 55.7 72.5
4 $\checkmark$ $\checkmark$ $\checkmark$ 68.5 96.5 56.8 73.5
@@
We evaluate the effectiveness of our fusion approach. As shown in entries #2 and #3 in Tab. 4, our frame-event fusion framework outperforms using either modality alone, demonstrating that our framework can effectively leverage the strengths of both modalities. The results from entry #4 reveal that concatenating the two modal images before feature extraction leads to performance degradation. The underlying reason may be that using independent parameters helps the model distinguish data types, enabling better learning of modality-specific features. In entry #5, we investigate an alternative approach of combining image features with event features via addition. This results in lower performance compared to our default method. Entry #6 demonstrates that modality-specific encoding more effectively differentiates modal features. Tracking experiments (entries #7, #8, and #9) confirm the necessity of our Decoupled Tracking Prediction With Similarity Fusion. Without this design, performance degrades to varying degrees. For more experiments, please refer to Appendix.
@@
We compare the tracking results of SpikeFET-Base with the state-of-the-art tracker shown in Fig. 2. Despite facing challenges such as exposure and low light conditions, our tracker achieves robustness and excellent tracking accuracy.

### v2

@@ Abstract
The integration of image and event streams offers a promising approach for achieving robust visual object tracking in complex environments. However, current fusion methods achieve high performance at the cost of significant computational overhead and struggle to efficiently extract the sparse, asynchronous information from event streams, failing to leverage the energy-efficient advantages of event-driven spiking paradigms. To address this challenge, we propose the first fully Spiking Frame-Event Tracking framework called SpikeFET. This network achieves synergistic integration of convolutional local feature extraction and Transformer-based global modeling within the spiking paradigm, effectively fusing frame and event data. To overcome the degradation of translation invariance caused by convolutional padding, we introduce a Random Patchwork Module (RPM) that eliminates positional bias through randomized spatial reorganization and learnable type encoding while preserving residual structures. Furthermore, we propose a Spatial-Temporal Regularization (STR) strategy that overcomes similarity metric degradation from asymmetric features by enforcing spatio-temporal consistency among temporal template features in latent space. Extensive experiments across multiple benchmarks demonstrate that the proposed framework achieves superior tracking accuracy over existing methods while significantly reducing power consumption, attaining an optimal balance between performance and efficiency.
@@ 3.1 Spiking Frame-Event Tracking Network
$\textbf{U}^{\prime}=\textbf{U}+\text{SepSpikeConv}(\textbf{U})$ (1)
@@
$\textbf{U}^{\prime}=\textbf{U}+\text{SepSpikeConv}(\textbf{U})$ (6)
@@
where $\text{Linear}(\cdot)$ denotes a linear transformation, $\gamma$ is the channel expansion factor (zhang2023rethinking, ), $\text{CSWinSSA}(\cdot)$ represents the CSWinSSA operator, and $\text{SSA}(\cdot)$ is the spiking self-attention operator within the CSWinSSA operator. To address the challenge of large values generated by matrix multiplication, a scaling factor (scale) is introduced to regulate the magnitude of the results. Further details of the CSWinSSA operator are provided in the Appendix E.
@@
In object tracking tasks, the response maps output by the network play a critical role in target localization. To achieve sufficient fusion of image and event modalities, we propose enhancing feature alignment by constraining the similarity between the response maps of the two modalities, implemented via a weighted focal loss function $\mathcal{L}_{\text{GWF}}$ (law2018cornernet, ) (see the Appendix C.1 for more details). The final loss function is expressed as: $\mathcal{L}_{\text{Res}}=\mathcal{L}_{\text{GWF}}\left(\mathbf{R}_{F}/\tau,\mathbf{R}_{E}/\tau\right)$, where $\mathbf{R}_{F}$ and $\mathbf{R}_{E}$ are the response maps of two modes, $\tau$ is the temperature coefficient, and the empirical setting is 2.
@@ 3.4 Training and Inference
where $\lambda_{\text{iou}}=2$ and $\lambda_{\text{L1}}=5$ are regularization parameters as in (yan2021learning, ). $\alpha=1$ and $\beta=0.5$ are hyperparameters used to balance the contributions of different loss functions. Ablation studies on hyperparameters are shown in Appendix  F.
@@
where $\lambda=0.5$ is a hyperparameter to balance the responses of different modalities. It should be noted that we do not employ a dynamic template update strategy. Instead, we consistently replicated the first frame to use two template frames (both derived from the first frame) and one search frame for inference. At the same time, we follow common practice by applying a Hanning window penalty to leverage positional priors in tracking (chen2021transformer, ; ye2022joint, ).
@@ 4.1 Implementation details
Our proposed SpikeFET was implemented with PyTorch 1.12 in Python 3.8 and trained on two NVIDIA RTX 4090 GPUs. The network input consists of a triplet image group comprising one search image and two distinct template images. When constructing training samples from randomly sampled video sequences, we expanded the search regions and template regions to 4× and 2× their original bounding box sizes, respectively, followed by resizing them to 256×256 and 128×128. Common data augmentation techniques such as horizontal flipping and brightness jittering were applied to the image sets. The model was trained using the AdamW optimizer. The optimizer has a cosine annealing scheduling over 50 epochs, where each epoch contained 60,000 image triplets. Please refer to the Appendix  D for more details.
Similar to current mainstream multi-modal object tracking frameworks (such as VIPT (zhu2023visual, ), SDSTrack (hou2024sdstrack, ), TENet (shao2025tenet, )), which commonly adopt OSTrack (ye2022joint, ) frame-datasets pre-trained models for parameter fine-tuning or knowledge distillation, we also follow their standard configuration process. Specifically, we adapt our SpikeFET into a single-modal SpikeET model, pre-trained on frame-datasets such as COCO (lin2014microsoft, ), LaSOT (fan2019lasot, ), TrackingNet (muller2018trackingnet, ), and GOT-10K (huang2019got, ) to fine-tune our SpikeFET network. For the single-modal SpikeET model, we remove one feature extraction branch and the spiking tracking head branch from SpikeFET, along with eliminating modality encoding. To balance speed and accuracy, we propose three variants with distinct architectures and parameters: SpikeFET-Base (Fig. 3) with larger parameter capacity, and SpikeFET-Tiny and SpikeET optimized for computational efficiency. Among these, only SpikeFET-Base undergoes frame-based dataset pre-training, while SpikeFET-Tiny and SpikeET adopt ImageNet-1K (deng2009imagenet, ) to pre-train the backbone, consistent with OSTrack (ye2022joint, ). To quantitatively evaluate the performance of tracking systems, we employ three core metrics: Area Under the Curve (AUC), Precision Rate (PR), and theoretical power consumption. For a detailed description of the theoretical power consumption calculation method, please refer to Appendix C.2.
4.2 Datasets
We evaluated our proposed SpikeFET using three large-scale frame-event single-object tracking datasets: FE108 (zhang2021object, ), VisEvent (wang2023visevent, ), and COESOT (tang2022revisiting, ). All three datasets were captured using the DAVIS346 camera, which features a spatial resolution of 346 × 260, a dynamic range of 120 dB, and a minimum latency of 20 $\upmu\text{s}$.
Specifically, The FE108 dataset (zhang2021object, ) contains 108 synchronized frame-event sequences captured in indoor environments, totaling approximately 1.5 hours of data. It covers 21 distinct target objects and is divided into 76 training sequences and 32 testing sequences. The target bounding box annotations were generated using a Vicon motion capture system. The VisEvent (wang2023visevent, ) dataset originally collected 820 frame-event pairs, divided into 500 training sequences and 320 testing sequences. Following (zhang2022spiking, ), we removed sequences with missing event data or misaligned timestamps, resulting in a refined dataset of 205 training sequences and 172 testing sequences. Multimodal networks and unimodal networks use 320 test sequences and 172 test sequences, respectively. COESOT (tang2022revisiting, ), as a larger-scale benchmark, contains 578K frame-event pairs divided into 827 training sequences and 527 testing sequences. Collected across diverse indoor and outdoor scenarios, it covers 90 target categories annotated with 17 challenging visual attributes. All target bounding boxes were manually annotated.
For fair comparison, we trained SpikeFET using the full training set of VisEvent and evaluated it on the test set containing 320 sequences, while SpikeET was trained on a pruned version of the VisEvent training set and tested on a subset of 172 sequences. Note that on SpikeFET, we only use raw event frames from the dataset on VisEvent, not event temporal frames.
@@ 4.3 Main Results
We evaluated the performance of our proposed SpikeFET-Tiny and SpikeFET-Base on several benchmarks, including FE108 (zhang2021object, ), VisEvent (wang2023visevent, ), and COESOT (tang2022revisiting, ). Meanwhile, for comparison with event-based trackers, we constructed a baseline model named SpikeET that using only event streams as input.
@@
Table 1: Comparison with state-of-the-art trackers on FE108 (zhang2021object, ), VisEvent (wang2023visevent, ), and COESOT (tang2022revisiting, ). We train and measure energy consumption on the VisEvent dataset. The best three results are shown in red, blue and green fonts. * indicates that frame-based trackers are extended to frame-event-based trackers through early fusion. $\dagger$ indicates models pre-trained on image-frame tracking datasets.
@@
ANN DiMP50* (bhat2019learning, ) - - 57.1 85.1 47.8 67.0 58.9 67.1
PrDiMP50* (danelljan2020probabilistic, ) - - 59.0 87.7 47.6 65.3 57.9 69.6
SiamRCNN* (voigtlaender2020siam, ) - - - - 49.9 65.9 60.9 71.0
TrDiMP50* (wang2021transformer, ) - - 60.3 91.2 - - 60.1 72.2
TransT50* (chen2021transformer, ) - - 63.9 93.0 47.4 65.0 60.5 72.4
ToMP101* (mayer2022transforming, ) - - 61.8 91.1 - - 59.9 67.2
FENet (zhang2021object, ) - 262.2 63.1 91.8 - - - -
@@
[…]
As shown in Tab. 1, compared to several state-of-the-art trackers, our SpikeFET demonstrates significant advantages across most datasets. On the FE108 (zhang2021object, ) dataset, SpikeFET-Tiny achieves an AUC score of 68.5%, matching the previous state-of-the-art HRMonTrack (zhu2023cross, ) while surpassing it by 0.3% in PR score and outperforming all other trackers by a substantial margin. Compared to the Tiny variant, the Base model further improves performance metrics, but there is a tendency towards overfitting. Notably, although the model size of the base model is comparable to the previous Base models, its theoretical energy consumption is significantly lower than equivalently scaled models. Additionally Compared with the latest SDSTrack (hou2024sdstrack, ) on ANN, our Tiny has improved its AUC score by 2.7% on the FE108 (zhang2021object, ) dataset, with parameters only one-third of SDSTrack (hou2024sdstrack, ), while reducing power consumption by 39 times.
@@
On the VisEvent (wang2023visevent, ) dataset, our Tiny model outperforms the majority of trackers. However, for the Base model, our SpikeFET performs comparatively worse than ViPT (zhu2023visual, ) and SDSTrack (hou2024sdstrack, ). We attribute this primarily to the fact that SpikeFET is particularly well-suited for handling datasets with challenging scenarios and large-scale datasets, such as FE108 (zhang2021object, ) and COESOT (tang2022revisiting, ). In such scenarios, the inherent advantages of event cameras become evident, allowing SpikeFET to demonstrate exceptional performance. In contrast, ViPT (zhu2023visual, ) and SDSTrack (hou2024sdstrack, ) excel in less challenging scenarios, such as VisEvent (wang2023visevent, ).
At the same time, as shown in Tab. 1, it can be observed that SpikeFET-Base achieved increases of 0.2%, 2.2%, and 4.5% in AUC metrics compared to SpikeFET-Tiny on the FE108 (zhang2021object, ), VisEvent (wang2023visevent, ), and COESOT (tang2022revisiting, ) datasets (arranged in ascending order of data volume), with the performance gap progressively widening. Tab. 2 reveals that on FE108 (zhang2021object, )—a dataset featuring numerous challenging scenarios (e.g., overexposure, low illumination)—SpikeFET’s performance is fully leveraged: its AUC is only 4.1% lower than SpikeFET-Tiny, while outperforming all event-based tracking algorithms. In contrast, it underperforms by 17.4% on VisEvent (wang2023visevent, ). This further illustrates the superiority of SpikeFET in challenging scenarios and complex datasets.
@@
ANN DiMP50 (bhat2019learning, ) - 256.37 - - 31.5 44.2
PrDiMP50 (danelljan2020probabilistic, ) - 258.37 - - 32.2 46.9
ATOM (danelljan2019atom, ) - 30.199 - - 28.6 47.4
SiamRPN (li2018high, ) - 203.88 - - 24.7 38.4
@@ 4.4 Ablation Study
We analyze the effects of the proposed RPM, $\text{E}_{\text{t}}$, and STR. Our baseline model, similar to CEUTrack (tang2022revisiting, ), processes six input images from both frame and event modalities with separate feature extraction. These features are concatenated before entering the fusion network, then separated prior to the prediction head and processed through decoupled tracking heads. As shown in Tab. 4, the baseline equipped with RPM (#3) achieves significant improvements. On FE108 (zhang2021object, ) with numerous extreme scenarios, RPM substantially increases AUC by 34.7% and PR by 49.9% over the baseline, while on VisEvent (wang2023visevent, ), it improves AUC by 10.4% and PR by 11.9%. More specifically, illustrated in Fig. 6, without RPM method, the peak response in the output image fails to precisely indicate the target position, exhibiting a significant deviation. In contrast, when RPM is employed, the target area is accurately highlighted with high response activation. Quantitative analysis confirms that our model achieves enhanced tracking accuracy and robustness through RPM.
Furthermore, we conducted further experiments using one template frame and one search frame per modality, totaling four inputs. Without RPM, the six-input configuration (#baseline) yields only marginally higher performance than the four-input setup (#1) on the FE108 (zhang2021object, ) dataset. Similarly, when RPM is employed, the six-input configuration (#3) also shows only a slight performance advantage over the four-input configuration (#2) on the FE108 (zhang2021object, ) dataset. This conclusively demonstrates that the primary innovation driving baseline improvement is RPM’s effective mitigation of padding-induced degradation in translation invariance. In contrast, utilizing additional data provides minimal performance gains relative to RPM’s contribution. Additionally, the six-input setting allows for the further incorporation of STR.
Further additions of $\text{E}_{\text{t}}$ (#4) and STR (#5) to #3 demonstrate effective enhancements, particularly with STR improving FE108 (zhang2021object, ) AUC by 0.6%. This proves that while retaining the advantage of RPM in dynamically reorganizing spatial distributions, the STR strategy effectively suppresses matching deviations caused by feature spatial misalignment, significantly improving model performance. Finally, integrating both $\text{E}_{\text{t}}$ and STR with RPM (#6) yields greater gains: 1.7% AUC and 1.5% PR improvements on FE108 (zhang2021object, ), along with 1.3% AUC and 1.1% PR improvements on VisEvent (wang2023visevent, ).
Table 3: Ablation studies on RPM, $\text{E}_{\text{t}}$, and STR (Random Patchwork Module, Type Encodings, Spatio-Temporal Regularization).
# Inputs RPM $\text{E}_{\text{t}}$ STR FE108 VisEvent
@@
Baseline 6 $\times$ $\times$ $\times$ 32.1 45.0 45.1 60.5
1 4 $\times$ $\times$ $\times$ 30.8 43.9 44.3 59.2
2 4 $\checkmark$ $\times$ $\times$ 66.1 93.9 54.5 71.6
3 6 $\checkmark$ $\times$ $\times$ 66.8 94.9 55.5 72.4
4 6 $\checkmark$ $\checkmark$ $\times$ 66.8 95.0 55.7 72.3
5 6 $\checkmark$ $\times$ $\checkmark$ 67.4 95.8 55.7 72.5
6 6 $\checkmark$ $\checkmark$ $\checkmark$ 68.5 96.5 56.8 73.5
@@
We evaluate the effectiveness of our fusion approach. As shown in entries #2 and #3 in Tab. 4, our frame-event fusion framework outperforms using either modality alone, demonstrating that our framework can effectively leverage the strengths of both modalities. The results from entry #4 reveal that concatenating the two modal images before feature extraction leads to performance degradation. The underlying reason may be that using independent parameters helps the model distinguish data types, enabling better learning of modality-specific features. In entry #5, we investigate an alternative approach of combining image features with event features via addition. This results in lower performance compared to our default method. Entry #6 demonstrates that modality-specific encoding more effectively differentiates modal features. Tracking experiments (entries #7, #8, and #9) confirm the necessity of our Decoupled Tracking Prediction With Similarity Fusion. Without this design, performance degrades to varying degrees. For more experiments, please refer to Appendix. F
@@
We compare the tracking results of SpikeFET-Base with the state-of-the-art tracker shown in Fig. 2. Despite facing challenges such as exposure and low light conditions, our tracker achieves robustness and excellent tracking accuracy. Additional visualization results are provided in the Appendix. A.